The perimeter of a rectangle is 42 ft. The length is 3ft longer than the width. Find the dimensions. Write a system of linear equation and solve the resulting system. Let x be the length and y be the width.
Answer by linh117(31)

x=3y
y
p=2(l+w)
42=2(3y+y)
42=8y
y=5.25
x=3*5.25
x=15.75